### v1

# Batch Bayesian Optimization for High-Dimensional Experimental Design: Simulation and Visualization  [^1] A footnote for the title

Abstract
[…]
keywords:
Batch Bayesian Optimization, Gaussian Process Regression, High Dimensions, Noise Effect, Synthetic Data, Batch-picking Methods
[…]
IR,NMR,UV
{tocentry}
Some journals require a graphical entry for the Table of Contents. This should be laid out “print ready” so that the sizing of the text is correct.
Inside the tocentry environment, the font used is Helvetica 8 pt, as required by Journal of the American Chemical Society.
The surrounding frame is 9 cm by 3.5 cm, which is the maximum permitted for Journal of the American Chemical Society graphical table of content entries. The box will not resize if the content is too big: instead it will overflow the edge of the box.
This box and the associated title will always be printed on a separate page at the end of the document.
1 Introduction
Optimizing from a rough process towards a fine-tuned set of process parameters has been the hallmark of success throughout human history. We can think about the shaping of tools during the stone age, the choice of the copper-tin composition in the bronze age, and the carbon content, furnace temperatures, and quenching conditions in the iron age. As our engineering capabilities have advanced throughout the industrial and the modern age, the number of parameters that can be changed during engineering processes have increased. Accelerating the optimization of engineering processes is critical to enable researchers and engineers to allocate more time to infer essential physical and chemical insights from experiments, thus facilitating faster scientific and industrial progress.
BO has recently emerged as a leading method for the efficient, sequential optimization of black-box functions that are costly to evaluate, 1, 2, 3, 4, 5 such as experimental work involving varying levels of automation.6 In BO, a surrogate model is constructed based on the cumulative data at each optimization iteration, and the next point(s) in the domain to test are selected based on the surrogate model according to a user-chosen acquisition policy. The choice of acquisition function and its exploration hyperparameter determines the balance between the exploitation of areas that have previously performed well and the exploration of the unknown regions. While the concept of BO is simple, i.e. to find the input that maximizes the black-box function, many subtleties of BO can befuddle new users.
First, BO operation is highly affected by its selected parameters. Gaussian process regression (GPR)7, 8 is the most popular BO surrogate model because the uncertainty of the model is automatically produced along with the posterior mean. The choice of GPR kernel and its hyperparameters, amplitude and lengthscales, represents the user’s beliefs on what the objective function looks like. Additionally, the GPR noise variance setting should reflect the noise level in the experiment. However, GPR is frequently used for materials science optimization problems that are so novel that the prior knowledge on the properties of the underlying objective function – e.g., its range, the domain landscape, noise level, or the number of optima – are not known beforehand.
The choice of acquisition function also presents another confusion to new users. Various acquisition functions have been presented in literature. Most BO literature, especially papers that are more mathematical and theoretical, use expected improvement (EI).9, 10 In contrast, the upper confidence bound (UCB) 11, 12 is simple and intuitive and is shown to be effective in experimental problems.13, 14 Users often do not know which one to use or how to set the hyperparameter that controls the tradeoff between exploitation and exploration.
Second, material science optimization often depends on many materials components and processing conditions, i.e., they are high-dimensional problems. Since the input parameters may not be independent of each other, it is pertinent to optimize them simultaneously to understand the interactions and dependencies better. With the increasing adaptation of automated/autonomous experiments, the correct implementation of high-dimensional BO algorithms becomes especially critical. However, up until recently, most BO implementations in materials science have only been discussed in low-dimensional, specifically, two- to three-dimensional BO. To our knowledge, the highest-dimensional published BO in experimental materials science work is in five dimensions 15, 16, and there are only a few works in four dimensions 17, 18, 19.
Due to the high dimensionality, the optimization progression is challenging to visualize; hence, evaluating the algorithm-guided experiment loops’ success (or failure) is not easy. A widely used approach in the literature is to use the model’s optimum value for the objective function as the optimization metric. However, this single metric for the BO is not capable of describing the results of the optimization campaign in detail. Furthermore, this presumes that the inputs that produce the optimized objective value of the model are the correct ones to produce the ground truth (GT) optimum, an assumption not yet verified.
Two large differences exist between experimental and simulation work: (1) experiments are often performed in batches, i.e., processing multiple samples at once, to save materials cost or time, and (2) experimental data contain aleatoric uncertainties manifesting as noise. Most BO work are performed for sequential optimization, i.e., picking one sampling condition for the next experiment based on the highest acquisition function value. When performing batch BO, after selecting the first next point at the highest acquisition function value, various selection strategies, categorized as penalizing, exploratory, or stochastic,20, 21, 22, 23 have been proposed to address how to select the remaining points so that they are meaningfully different from the first point and each other. This is challenging because we do not gain any new information between selecting the first and the rest of the batch. Noise is inevitable in experiments and may have a substantial impact to the outcomes in high-dimensional optimization tasks.24, 25, 26, 27, 28 Unfortunately, most of the popular BO algorithm choices implemented in open BO code repositories have been developed and benchmarked in scenarios with negligible experimental noise. Therefore, it is unclear if the same choices of algorithm, acquisition function, or hyperparameter perform well in noisy, high-dimensional experimental tasks encountered in materials science. A systematic evaluation of the effects of noise on optimization outcomes is needed. In simulated BO studies, noise is commonly added based on the proportion of the GT objective value (y) maximum. Depending on the shape of the objective function, this might not reflect the general signal-to-noise level in experiments.
Finally, before implementing BO in an experimental campaign, it is good practice to test BO in a simulation environment using synthetic data, for pedagogical and troubleshooting purposes. In this work, we develop a framework to visualize BO step by step, as an evaluation tool for simulation environments and as a debugging tool for experiments. We showcase an example of simulated data with increasing noise, evaluating optimization strategies as a function of noise magnitude. In our demonstration, we implement batch BO using the Emukit package to find the optimum to two different types of 6-dimensional (6D) objective functions: a needle-in-a-haystack function and a function with a local optimum nearly degenerate with the global optimum. 29, 30
2 Method
2.1 Synthetic objective functions for benchmarking and analysis of Bayesian optimization
[…]
font=footnotesize Refer to caption
Figure 1: : Visualization of the 3D representation of the (a) Ackley function ground truth, where x1,x2 variables are projected in 3D representation, and (b) Hartmann function ground truth, where x1,x3 variables are projected in 3D representation. The global maximum is labeled as the ‘GT max’ (cyan cross for Ackley, black cross for Hartmann) both functions and Hartmann function 2nd maxima labeled as the ’2nd True Max’ (red star).
We consider two synthetic test functions: (1) an Ackley function31
$A({\bf X})=20\left(\exp\left(-0.2\sqrt{\frac{1}{d}\sum_{i=1}^{d}x_{i}^{2}}\right)-1\right)+\exp\left(\frac{1}{d}\sum_{i=1}^{d}\cos(2\pi x_{i})\right)-\exp(1)$ (1)
where the dimension $d$ = 6, the range of $A({\bf X})$ is $[-22.3,0]$ whereas the domain ${\bf X}$ is a 6D hypercube spanning $[-32.768,32.768]$ along each dimension. Note that compared to Ref. 31, we changed the sign of the Ackley function so that the optimization problem is a maximization problem. The Ackley function, illustrated in Fig.1(a)(only x1,x2 pair variables projected on Ackley function GT) and Fig. S1(all pairs of variables projected on Ackley function GT), has one sharp global maximum $A({\bf X}_{\rm max})=A({\bf 0})=0$ at the origin. The Ackley function thus represents a needle-in-the-haystack-type problem. Functions of this type are also referred to as heterogeneous functions,32, 33 where the convex region around the maximum occupies a minuscule fraction of the parameter space. Specifically, greater than 99.99% of the hypervolume of the domain has $f({\bf X})<-15$, or, equivalently and more precisely, when randomly sampling ${\bf X}$, the probability of finding $f({\bf X})>-15$ is less than $4\times 10^{-5}$.
(2) a Hartmann function31
$H({\bf X})=\frac{1}{1.94}\left[2.58+\sum_{i=1}^{4}\alpha_{i}\exp\left(-\sum_{j=1}^{d}A_{ij}(x_{j}-P_{ij})^{2}\right)\right]$ (2)
where $d$ = 6, $\alpha=(1.0,1.2,3.0,3.2)^{T}$ , $A_{ij}$ and $P_{ij}$ are $4\times 6$ matrices defined in Supplementary section 1. The range of $H({\bf X})$ is $[0,3.32237]$ whereas the domain ${\bf X}$ is a 6D unit hypercube spanning from $[0,1]$ along each dimension. We visualize the Hartmann function in Fig.1(b)(only x1,x3 variables pair projected on Hartamnn function GT) and Fig. S2 ( all pairs of variables projected on Hartmann function GT). The Hartmann function has one global maximum $H({\bf X}_{\max})=3.32237$, located at ${\bf X}_{\max}=(0.20169,0.150011,0.476874,$ 0.275332,0.311652,0.6573)0.275332,0.311652,0.6573)0.275332 , 0.311652 , 0.6573 ). A second local maximum $H({\bf X}_{\max,2})=3.20452$ is found at ${\bf X}_{\max,2}=(0.20169,0.150011,0.476874,$ 0.275332,0.311652,0.6573)0.275332,0.311652,0.6573)0.275332 , 0.311652 , 0.6573 ). The gradients near the maximum of the Hartmann function change much more gradually compared to the Ackley function, and we say that the Hartmann function is non-heterogeneous. The Hartmann function represents a shallow maximum and a landscape where optimization can easily get trapped in a local maximum far away from the global maximum.
2.2 Main parameters of Bayesian optimization considered during benchmark and analysis
The most utilized and investigated surrogate model for BO is GPR. The posterior distribution of a GPR model at a point ${\bf X}$ is $f({\bf X})\sim N({\bf m},{\bf K})$ where $N$ refers to the normal distribution, ${\bf m}$ the mean vector, and ${\bf K}$ the covariance matrix. Given a set of observed data points ${\bf D}^{n}=\left\{{\bf X}^{i},y^{i}\right\}_{i=1}^{n}$, $i$ being the sample index and $n$ being the number of samples that has been evaluated so far, for a new point ${\bf X}^{i+1}$, the posterior predictive distribution can be computed as, $y^{n+1}\sim N(\mu_{\bf D}({\bf X}^{n+1}),(\sigma_{\bf D})^{2}({\bf X}^{n+1}))$, where $\mu_{\bf D}$ is the predicted mean and $(\sigma_{\bf D})^{2}$ is the variance, computed from ${\bf D}^{n}$.
The first step in building a GPR model in BO is choosing a utility function $u({\bf D})$, where a higher $u$ indicates that the dataset ${\bf D}$ is of higher quality and better able to specify where the global maximum is.34, 35 The most naive choice for the utility function is $u({\bf D})={\rm Max}(y)$, however, this utility function can improve in the presence of noise which is not meaningful. Instead, we use the utility function
$u({\bf D})={\rm Max}(\mu_{\bf D}({\bf X}))$ (3)
where, $\mu_{\bf D}$ is the predicted mean computed using ${\bf D}$, and the maximum (${\rm Max}$) is considered over all ${\bf X}$ in ${\bf D}$. Eq. 3 approaches $u({\bf D})={\rm Max}(y)$ in the absence of noise, but will be a better utility function in the presence of noise.
We consider the EI and UCB acquisition policies. The EI acquisition function is expressed explicitly in terms of the predicted mean $\mu_{\bf D}$ and variance $(\sigma_{\bf D})^{2}$ as
${\rm EI}({\bf X}|\mu_{\bf D},(\sigma_{\bf D})^{2})=(\mu_{\bf D}({\bf X})-u({\bf D})-\xi)\phi(Z({\bf D}))+\sigma_{\bf D}({\bf X})\varphi(Z({\bf D}))$ (4)
$Z({\bf D})=\frac{\mu_{\bf D}({\bf X})-u({\bf D})-\xi}{\sigma_{\bf D}({\bf X})}$ (5)
where $\phi(Z)$ indicates cumulative distribution function (CDF) and $\varphi(Z)$ indicates probability distribution function (PDF) of the GPR surrogate model. $\xi>0$ is an exploration hyperparameter; the larger the $\xi$, the more aggressive the exploration. The UCB acquisition function is defined by,
$\rm UCB({\bf X}|\mu_{\bf D},(\sigma_{\bf D})^{2})=\rm\mu_{\bf D}({\bf X})+\beta(\sigma_{\bf D}({\bf X}))^{2},\beta>0$ (6)
where $\beta$ represents an exploration hyperparameter. We selected UCB and EI because they are used widely, have been implemented in multiple BO code packages, and can tune exploration and exploitation easily via hyperparameters $\beta$ and $\xi$, respectively. In this work, we benchmark both acquisition functions with a range of exploration/exploitation hyperparameter values to evaluate how they affect BO learning outcome.
To mimic experiments in fabricating/testing of multiple samples in a round, we perform batch BO with a batch size of four. We compare three batch-picking methods: Local Penalization (LP),36 Kriging Believer (KB),37 and Constant Liar (CL),38, 39 representing penalizing, exploratory, stochastic strategies and commonly implemented in open BO packages. 40 Although BO was created as a sequential optimization tool, employing batch evaluation is critical to leverage parallelization.
2.3 Benchmark Approach
[…]
font=footnotesize Refer to caption
Figure 2: The workflow in the batch BO benchmarking: (a) LHS of the 6D input variable space to pick the starting points for the BO, (b) evaluations of the analytical test function at the selected points with an option to include noise, (c) the surrogate GPR model training at each iteration of the BO learning, (d) picking a batch of input points for the next iteration. The whole BO learning runs 50 iterations to generate the (e) X (top) and y (bottom) learning curves of the benchmark criteria selected for this work, tracking the distance of the surrogate model optimum point to the true optimum location and the value of the surrogate model optimum, respectively. This whole process is repeated for 99 different LHS initial samplings to collect statisticsBO.
We benchmark BO with the two test objective functions (A($\bf X$) and H($\bf X$)), two acquisition functions (EI and UCB) with respective hyperparameters, as well as three batch-picking methods (LP, KB, and CL), without or with different levels of noise according to the workflow illustrated in Fig. 2. The workflow involves initial sampling (Fig. 2(a)), repeating the BO iterations which include evaluating the objective values at each sampled point (Fig. 2(b)), model training (Fig. 2(c)), and suggesting the input points to evaluate for the next batch of points (Fig. 2(d)), and evaluating the learning curves after the end of BO, in our case after 50 iterations (Fig. 2(e)). BO settings, apart from the benchmark variables, are kept the same across the benchmark simulations, as detailed next.
As illustrated in Fig. 2(a), the BO was initialized with 24 ${\bf X}$ points using the LHS method. Figs. 2(b)-(d) show how noise (if any) is added to the objective function values at the sampled points, the BO model is trained based on the accumulated data points, and a batch of ${\bf X}$ is chosen based on the acquisition function and batch-picking method. We considered BO in a 6D space with 4 points per batch and analyzed the progress through learning curves after 50 iterations as illustrated in Fig. 2(e). Internally, both the input ${\bf X}$ and their corresponding $y$ values are normalized to the respective ranges of the ground-truth functions, i.e., the input and output variable values are within the unit hypercube. In our figures, we present results scaled back to their actual range.
Noise is generated according to,
$y^{i}=\rm f({\bf X}^{i})+\epsilon^{i},\epsilon^{i}\sim\textit{N}(0,(\sigma^{i})^{2})$ (7)
We introduce noise in two distinct methods to analyze its impact on BO. Following the common practice in the literature 25, noise was added as a percentage of the maximum GT value,
$\sigma^{i}={\rm Max}(y_{GT})\times\textrm{proportion\; of\; noise}$ (8)
In experiments, the GT is unknown, and thus the noise level is typically characterized as signal-to-noise ratio (SNR). To represent SNR in the synthetic data, we argue that the kernel amplitude in the noiseless case represents the general level of the signal better than its global maximum value, and hence noise should be added as a percentage of the noiseless kernel amplitude,
$\rm\sigma^{i}=\textrm{(kernel\; amplitude)}_{\epsilon=0}\times{\rm proportion\;of\;noise}$ (9)
The results of these two ways to incorporate noise are compared.
Model training (Fig. 2(c)) involves training a GPR surrogate model at each iteration of BO. In each scenario, the ARDMatern52 is used as the kernel function. Automatic relevance determination (ARD) kernels assume a different length scale for each dimension and tune them separately. 41, 42 In a previous benchmark by Liang et al., the ARD kernels performed better than the non-ARD ones.1 The ARD Matern kernel function is:
$\rm K_{M}(x1,x2)=\sigma^{2}\frac{2^{1-\nu}}{\tau(\nu)}\left(\frac{\sqrt{2\nu}|x1-x2|}{L}\right)^{\nu}K_{\nu}\left(\frac{\sqrt{2\nu}|x1-x2|}{L}\right)$ (10)
where $\sigma^{2}$ is the amplitude parameter, L is a length scale parameter, $\tau(.)$ is the gamma function, $\rm K_{\nu}$ is a modified Bessel function of the second kind and $\nu$=5/2 refers to the smoothness of the function (lower value means more smooth).5, 43 The kernel amplitude and length scales are hyperparameters that must be correctly tuned to construct a realistic GPR surrogate model. Kernel amplitude represents the value range in the surrogate model predictions, and length scale represents the correlation between two points in the input space $\bf X$ of the surrogate model. Additionally, GPR model has a Gaussian noise variance (GNV) hyperparameter, representing the uncertainty associated with each observation and, when applied to experiments, reflecting the noise level in the data. Therefore, the GPR surrogate model may overly smooth the observed data if the noise variance is too high, whereas the model may overfit the noisy observations if the noise variance is too low. Both cases lead to suboptimal decisions. All three hyperparameters are autotuned in the simulation and tracked with iterations.
The next four points to sample are suggested according to the acquisition function and batch-picking method (Fig. 2(d)). We test LP, KB, and CL as batch-picking methods for both Ackley and Hartmann functions with both EI and UCB as the acquisition functions with optimized hyperparameters based on performance metrics. We find LP outperforms KB and CL for both functions independent of the choice of acquisition function. For details see Supporting Information appedix section D. For the rest of the paper, LP is used as the batch-picking method unless explicitly specified.
We always consider 99 different initial LHS samplings which we indicate with a subscript index $k$ and 50 iterations which we indicate with a superscript $i$. When an index is omitted, it implies the collection of all ${\bf X}$ under consideration. We define ${{\bf X}_{k}^{i}}^{*}$ as the ${\bf X}$ associated with ${\rm Max}(\mu_{\rm D})$ up to the current iteration, the formal definition is
$\mu_{\rm D}({{\bf X}_{k}^{i}}^{*})={\rm Max}_{j=1}^{24+4i}(\mu_{\rm D}({\bf X}_{k}^{j}))$ (11)
where the maximum is taken over all 24 evaluations from the LHS sampling and the $4i$ evaluations in the first $i$ iterations of the BO. $\mu_{\rm D}({{\bf X}}^{*})$ is the same as ${\rm Max}(\mu_{\rm D}({\bf X})$ defined in Eq. 3.
2.4 Metrics for Performance Evaluation During Benchmarking
To characterize the optimization of the BO results, we compute instantaneous regret (IR) for ${\bf X}$ and $y$: ${\rm IR}({\bf X}_{k})=||{{\bf X}_{k}^{50}}^{*}-{\bf X}_{\rm max}||$, and ${\rm IR}(y_{k})=|\mu_{\rm D}({{\bf X}_{k}^{50}}^{*})-y_{\rm max}|$, and then average the IR over all 99 LHS samplings to establish statistical variation:
$\langle{\rm IR}({\bf X})\rangle=\frac{\sum_{k=1}^{99}||{{\bf X}_{k}^{50}}^{*}-{\bf X}_{\rm max}||}{99}$ (12)
and
$\langle{\rm IR}(y)\rangle=\frac{\sum_{k=1}^{99}|\mu_{\rm D}({{\bf X}_{k}^{50}}^{*})-y_{\rm max}|}{99}.$ (13)
To quantify the convergence rate, average cumulative regrets in X and y are calculated:
$\langle{\rm CR}({\bf X})\rangle=\frac{\sum_{k=1}^{99}\sum_{i=1}^{50}||{{\bf X}_{k}^{i}}^{*}-{\bf X}_{\rm max}||}{99}$ (14)
and
$\langle{\rm CR}(y)\rangle=\frac{\sum_{k=1}^{99}\sum_{i=1}^{50}|\mu_{D}({{\bf X}^{i}_{k}}^{*})-y_{\rm max}|}{99}.$ (15)
For all four metrics, smaller values indicate a better performance.
3 Results and Discussion
3.1 Noise-free
3.1.1 Ackley
Noise-free optimizations are investigated first to determine the ideal performance of BO on our optimization tasks, and to illustrate the convergence of the BO in terms of our performance evaluation metrics. Fig. 3 shows the BO of the Ackley function using the UCB acquisition policy. Fig. 3(a) shows $||{\bf X}_{k}^{i}-{\bf X}_{\rm max}||$, the deviation of the current optimal location from the GT maximum location (${\bf X}_{\rm max}$) as a function of iteration index $i$ for all 99 LHS samplings. Since the initial pick is random, none of the 99 samplings yield initial vectors close to ${\bf X}_{\rm max}$. Thus, $||{\bf X}^{1}-{\bf X}_{\rm max}||>15$ which is not surprising since the search space is a large hypercube measuring $[-32.768,32.768]^{6}$. After 10 iterations significant progress towards the optimum is made, $||{\bf X}^{10}-{\bf X}_{\rm max}||<10$ for all 99 samplings. At the last iteration ${\rm IR}({\bf X}_{k})=||{\bf X}_{k}^{50}-{\bf X}_{\rm max}||<1$ and the estimated optimal inputs for all 99 samplings are very close to the GT maximum.
[…]
font=footnotesize Refer to caption
Figure 3: BO results using UCB with $\beta$ = 1 on noise-free Ackley function. Learning curve in (a) $||{\bf X}^{*}-{\bf X}_{\rm max}||$ and (b) $\mu_{\rm D}({\bf X}^{*})$ for all 99 LHS initial starts. The 25th percentile (green triangle), 50th percentile (red circle) and 75th percentile (blue triangle) regions are highlighted to exemplify poor, median, and good LHS BO models, respectively. (c) Visualization of the 3D representation (x1 vs. x2) for the 50th percentile BO model, showing how BO iterations zero in on ${\bf X}_{\rm max}$ (cyan cross). Light blue circles are the 24 initial LHS selections and the pink to dark red points are training points (progressively darker). (d) Zoomed-in 2D heat map for variables x1 and x2 within the range [-2, 2] near ${\bf X}_{\rm max}$. (e) Parity plot of the 50th percentile BO model prediction vs. GT values for all 224 sampled points (LHS + 50 iteration at a batch size of 4).
After 50 iterations, the 99 samplings are ranked based on ${\rm IR}({\bf X})$ from 1 to 99 percentile (worse to best). In Fig. 3, we highlight the evolution of the sampling that is ranked 25th (poor outcome), 50th (median outcome), and 75th (good outcome) in green, red, and blue, respectively. The percentile ranking is made based solely on the result after the 50th iteration, and therefore the order can change throughout the iterations. Indeed, upon careful inspection, we observe that the 75th ranked sampling is further away from the GT compared to the 50th ranked sampling for the first six iterations.
Fig. 3(b) shows the $y$ value estimated by the surrogate GPR model, $\mu_{\rm D}({\bf X}^{*})$, as a function of BO iterations. At iteration 1, all $y$-value estimates are below $-15$, whereas around iteration 20 estimates have improved to be above $-5$, and at the final iteration the $y$-values are all close to 0, the maximum $A({\bf X})$ value. Fig. 3(c) shows a 3D representation of the surrogate model and scatter plot of the inputs of the 50th percentile sampling on two of the six ${\bf X}$ dimensions. Along the input range from -32.768 to 32.768, the objective function ranges from - 22.3 to 0. Each dot indicates an input point; blue dots indicate initial LHS points and the red dots from light to dark represents the learning progression. As iterations progress more points are sampled close to ${\bf X}_{\rm max}$, indicating that the BO exploits the optimum region. Fig. 3(d) shows the 2D heat map with focused domain range near maximum $A({\bf X})$ and reveals that most of the points for the Ackley function are sampled around ${\bf X}_{\rm max}$.
In addition to the analysis of performance metrics, the evolution of GPR hyperparameters is provided in Fig. S3. We observe that all length scales converged to similar values as expected from the rotational symmetry around the origin of the Ackley function. It should be noted that the exploration hyperparameter $\beta$ greatly affects BO convergence.34, 12 Table  S1 shows the different metrics, $\rm\langle IR({\bf X})\rangle$, $\rm\langle CR({\bf X})\rangle$, $\rm\langle IR(y)\rangle$, and $\rm\langle CR(y)\rangle$, all normalized to the ${\bf X},y$ ranges, for different $\beta$ values. We chose the $\beta$ that produces the smallest $\rm\langle IR(\bf X)\rangle$ which we determine as $\beta=1$, which also yields the smallest $\langle{\rm CR}({\bf X})\rangle$ and $\langle{\rm CR}(y)\rangle$.
Fig. 3(e) shows the parity plot for the 50th percentile sampling. The parity plot shows the GPR model posterior mean predictions as a function of their corresponding GT function values. The overall root mean square error (RMSE) of the posterior mean predictions on the training data is $\sim 4\times 10^{-4}$. Normalizing the inputs to the unit hypercube and the outputs to $[0,1]$, the RMSE is $\sim 1.8\times 10^{-5}$, reflecting the square root of the final value of the GNV hyperparameter of the model (Fig. S3(h)). Although our data does not contain noise, a non-zero value is needed for the GNV hyperparameter to ensure convergence. We note that when the GNV hyperparameter is set to be proportional to the mean squared variance of the unnormalized $y$ values, the GPR model deviates from the GT function as if noise is added, indicating the GNV is too large. Fig. S4 shows how the learning outcomes are affected by the GNV values. Judiciously chosen GNV value ensures small RMSE value, i.e., the GPR model reproduces the GT Ackley function very well when data are without noise.
[…]
font=footnotesize Refer to caption
Figure 4: (a)-(f) 3D representations of Ackley test function (UCB with $\beta$=1) at different iterations. Blue points are the initial LHS points. Light pink points are sampled at earlier iterations and the dark red points are sampled at later iterations according to the color bar.
To visualize the learning process even better, we provide Movie (1) showing the evolution of the median GPR model where the iteration progress is represented through the time evolution. Fig. 4 shows still snapshots of the evaluated points and the associated GPR model taken from Movie (1), clearly depicting how the model develops into something similar to the GT function as BO progresses and the number of data collected increases.
A concern with using the Ackley function for simulating BO is that its optimal point is located at the center of the domain, making it possible for the search to be accidentally expedited if using grid-sampling algorithm, due to the steepness of the maximum. Thus, we do not use the grid-sampling method in this work. In Fig. S5, we test our BO algorithm on an Ackley function with a maximum off the center at (x1, x2, x3, x4, x5, x6). The convergence of BO is similar to Fig. 3 and the maximum is correctly identified.
3.1.2 Hartmann
Fig. 5a shows $||{\bf X}_{k}^{i}-{\bf X}_{\rm max}||$ for the Hartmann test function using the UCB acquisition policy. Initially, the distance to the maximum $||{\bf X}^{1}_{k}-{\bf X}_{\rm max}||$ ranges from $0.2$ to $1.5$ in the unit hypercube. The distance to ${\bf X}_{\rm max}$ for the Hartmann function show relatively larger variance compared to the Ackley, which we attribute to those LHS sampling result in finding a optimum close to ${\bf X}_{\rm max,2}$ rather than ${\bf X}_{\rm max}$. In Fig. 5(a), for $\approx 75\%$ of the samplings, $||{\bf X}_{k}-{\bf X}_{\rm max}||$ converges to 0 while for the remaining $\approx 25\%$, $||{\bf X}_{k}-{\bf X}_{\rm max}||$ converges to $||{\bf X}_{\rm max,2}-{\bf X}_{\rm max}||\approx 1.1$. In some LHS samplings ${\bf X}_{k}$ is initially near ${\bf X}_{\rm max,2}$ but still ends up at ${\bf X}_{\rm max}$. Overall, because of the second maximum, it is significantly more challenging to find the GT max in the Hartmann landscape compared to the Ackley landscape.
[…]
font=footnotesize Refer to caption
Figure 5: BO results using UCB with $\beta$ = 1 on noise-free Hartmann function. Learning curve in (a) $||{\bf X}^{*}-{\bf X}_{\rm max}||$ and (b) $\mu_{\rm D}({\bf X}^{*})$ for 99 LHS BO models. The 25th percentile (green triangle), 50th percentile (red circle) and 75th percentile (blue triangle) regions are highlighted to exemplify poor, median, and good LHS BO models, respectively. (c) Visualization of the 3D representations for (x1 vs. x2) variables pair, for the 50th percentile BO model at the last iteration, showing how BO iterations zero in on ${\bf X}_{\rm max}$. Red circles are the 24 initial LHS selections and the blue to light green points are training points (blue progressively to green).(d) Parity plot of the 50th percentile BO model prediction vs. GT values for all 224 sampled points (LHS + 50 iteration at a batch size of 4).
Fig. 5b shows $\mu_{\rm D}({\bf X}^{*})$ as a function of iteration index for all 99 samplings. Initially, the values range from 0.5 to 3.0 (whereas the GT maximum is 3.32). The relative variance of $\mu_{\rm D}({\bf X}^{*})$ is much greater compared to the Ackley landscape. A number of LHS samplings find initial ${\bf X}$ which are already relatively close to $H({\bf X}_{\rm max})$. The reason is that the global maximum is wide and there is also a second maximum with almost as high function values, thus a much larger fraction of the domain yields values close to $H({\bf X}_{\rm max})$. After 10 iterations, all $\mu_{\rm D}({\bf X}^{*})$ are above 2.5 and then quickly converge to the GT maximum (3.32237) or the second ground truth maximum (3.20452).
The dependence of the BO results on the UCB exploration hyperparameter $\beta$ is shown in Table  S3. Similarly to Ackley, $\beta$ = 1 produces the lowest $\rm\langle IR(y)\rangle$ when using UCB policy. Comparing Table  S1 and  S3 for $\beta$ = 1 shows that ${\rm\langle IR}({\bf X})\rangle$ and $\rm\langle IR(y)\rangle$ are $\sim$ 210 times larger and $\sim$ 2 times smaller, respectively, for the Hartmann compared to the Ackley function. Our findings are in line with the benchmarking study of Liang et al.1
In Fig. 5(c), the ${\bf X}$ evaluated at each iteration of the 50th percentile (median outcome) LHS sampling are shown superimposed on the 3D representations of the GPR model. The red dots are LHS points and the blue-to-green dots represent BO learning progression. The projections onto the other pairs of input variables are shown Fig.  S6. The GPR model of the median sampling, hence all the percentile samplings above the median, accurately identify ${\bf X}_{\rm max}$ of the test function. The sampled points (blue to light green circles) converge on ${\bf X}_{\rm max}$ and no evaluated points appear near ${\bf X}_{\rm max,2}$ demonstrating that in this case, the BO model correctly identifies ${\bf X}_{\rm max}$. The parity plot in Fig. 5(d) compares the median BO model predictions with the GT data for all the evaluated points. RMSE value for this plot is $2.6\times 10^{-4})$. The normalized RMSE value is $7.8\times 10^{-5}$, which corresponds well to the converged square root value of the GNV hyperparameter (Supplementary Fig.  S7(h)).
Fig. S8 shows still snapshots of the BO iterations for the 50th percentile LHS sampling, projected onto the x1-x3 plane, whereas Movie (2) shows the entire iterative progression of the BO. Note that in Fig. 5(c) the 3D representation is of the GP model of the final iteration whereas in Fig. S8, the evolution of the GPR model is illustrated. Initially, the GPR model displays a very flat landscape without a peak and the $\mu_{\rm D}({\bf X}^{*})$ value is $<2$. As iterations progress, the GPR model develops a peak near the GT max ${\bf X}_{\rm max}$ (black cross). By the tenth iteration, the BO model correctly identifies the ${\bf X}_{\rm max}$. Note that there is no peak near the $2^{\rm nd}$ max ${\bf X}_{\rm max,2}$ (red star).
We also analyzed poor (25th percentile) and good (75th percentile) BO models based on the ranking of their final ${\rm IR}({\bf X})$ values (see Supplementary Movie (3) and (4), respectively). The poor GPR model has the optimum near ${\bf X}_{\rm max,2}$; on the contrary, the good model had the optimum exactly at the ${\bf X}_{\rm max}$. The latter also had a comparatively faster convergence than the median LHS optimized model as indicated in Fig. 5 (a) and (b).
3.1.3 Comparing the UCB and the EI acquisition functions
The UCB and EI acquisition policies are compared for the Ackley and the Hartmann function optimization in Table 1.
[…]
font=footnotesize Test Function Acquisition Policy $\rm\langle IR({\bf X})\rangle$ 1 ×$\times 10^{-2}$ $\rm\langle CR({\bf X})\rangle$ $\rm\langle IR(y)\rangle$ 1 ×$\times 10^{-2}$ $\rm\langle CR(y)\rangle$ Ackley UCB ($\beta$ = 1) 0.11 1.17 1.63 4.88 Ackley EI ($\xi$ = 0) 0.59 1.87 9.35 8.61 Hartmann UCB ($\beta$ = 1) 23.1 18.5 0.84 3.31 Hartmann EI ($\xi$ = 0) 33.0 18.9 0.91 3.10
Table 1: Comparison between UCB and EI acquisition policies with noise-free Ackley and Hartmann test functions. Exploration parameter settings ($\beta$ for UCB and $\xi$ for EI. respectively) are also listed.
It lists our four BO performance evaluation metrics that rely on instantaneous and cumulative regrets (${\rm\langle IR}({\bf X})\rangle$, ${\rm\langle CR}({\bf X})\rangle$, ${\rm\langle IR}(y)\rangle$, and ${\rm\langle CR}(y)\rangle$, the lower the value, the better the performance). The EI acquisition policy exploration hyperparameter $\xi=0$ was determined through optimization, similarly to the UCB acquisition function hyperparameter $\beta$. $\xi$ = 0 resulted in the lowest ${\rm\langle IR}({\bf X})\rangle$ on both test functions, as shown in Table  S2 and Table  S4. Independent of the metric or objective function, UCB yields lower regret values except for $\rm\langle CR(y)\rangle$ for Hartmann, where EI produces a slightly lower value (3.10) than UCB (3.31). The difference is observed in ${\rm\langle IR}({\bf X})\rangle$ for the Ackley function where the UCB value is more than 5 times lower. Figs. S9 and  S10 show the superior performance of UCB compared to EI visually via the evolution of $||{\bf X}^{*}-{\bf X}_{\rm max}||$ and $\mu_{\rm D}({\bf X}^{*})$, with the most significant difference for the Ackley test function. Overall, we conclude that UCB significantly outperforms the EI, especially in terms of the convergence of $\mu_{\rm D}({\bf X}^{*})$, which is also in line with Liang et al.1.
3.2 Noise
3.2.1 Comparing Utility functions in the presence of noise
Fig. 6a-d compares the learning curves using ${\rm Max}(y)$ (a, c) vs $\mu_{\rm D}({\bf X}^{*})$ (b, d) utility functions with 5% noise for both Ackley (a, b) and Hartmann (c, d) test functions. The slower convergence in Fig. 6 compared to Figs. 3 and 5 is due to the presence of noise in the data. The ${\rm Max}(y)$ results (left column) reveal that about 5 LHS samplings yield ${\rm Max}(y)>A({\bf X}_{\rm max})$ for Ackley, and the majority of LHS samplings produce ${\rm Max}(y)>H({\bf X}_{\rm max})$ for Hartmann. $y$ can exceed the GT maximum value because of the noise when evaluating ${\bf X}$ close to the maximum or second maximum. In such a case, the learning curve is tracking the outliers, i.e., data point with the largest noise value, and the problem is exacerbated at higher noise levels. In contrast, when $\mu_{\rm D}({\bf X}^{*})$ is used as the utility function in the BO, almost no LHS sampling result in $\mu_{\rm D}({\bf X}^{*})$ exceeding the GT maximum value. Furthermore, the $\mu_{\rm D}({\bf X}^{*})$ learning curve of the Ackley test function (Fig. 6b) reveals more clearly that the algorithm has not yet converged to the optimum compared to the ${\rm Max}(y)$ curve (Fig. 6a). Overall, we conclude that using $\mu_{\rm D}({\bf X}^{*})$ as the utility function for the learning curves is a more suitable measure of BO algorithm than Max(y), especially when noise is present.34 We have tracked $\mu_{\rm D}({\bf X}^{*})$ in all other learning curves in the paper.
[…]
font=footnotesize Refer to caption
Figure 6: Learning curves of (a) Max(y) and (b) $\mu_{\rm D}({\bf X}^{*})$ for Ackley test function and of (c) Max(y), and (d) $\mu_{\rm D}({\bf X}^{*})$ for Hartmann test function using EI as acquisition function ($\beta$ =1) with 5% noise level. The 25th percentile (green triangle), 50th percentile (red circle) and 75th percentile (blue triangle) BO models are highlighted to exemplify poor, median, and good LHS BO models, respectively. The yellow dash line represents the GT global maximum value of the test functions, $A({\bf X}_{\rm max})$ or $H({\bf X}_{\rm max})$.
3.2.2 Ackley
Fig. 7 (a), and (b), show the learning curves of the Ackley test function for noise levels of 2%, 5%, 7%, and 10%, from left to right, using EI as the acquisition policy. With 2% noise, the curves look similar compared to the no-noise case although at iteration 10, one LHS sampling does not satisfy $||{\bf X}^{*}-{\bf X}_{\rm max}||<10$ whereas all satisfied this criterion for the noiseless case. At 5% noise, we observe that many samplings do not reach the GT maximum but the 50th percentile sampling still reaches a value close to ${\bf X}_{\rm max}$. This indicates that a little more than half of the samplings find the global maximum location by the end of BO despite $\mu_{\rm D}({\bf X}^{*})$ is lower than $A({\bf X}_{\rm max})$.
[…]
font=footnotesize Refer to caption
Figure 7: Learning curve of $||{\bf X}^{*}-{\bf X}_{\rm max}||$ in top row (a) and $\mu_{\rm D}({\bf X}^{*})$ in the second row (b) for the Ackley function using LP as the batch-picking method and EI as the acquisition function for noise levels of 2% ($\xi$ = 0), 5% ($\xi$ = 0.05), 7% ($\xi$ = 0.1), and 10% ($\xi$ = 0.05) from left to right. For visualization, (c) 3D representations and (d) parity plots of the 50th percentile BO models for the same noise levels. In the 3D representations, x1 and x2 are projected with 224 learning points, including 24 initial LHS points (blue) and 50 iterations (from light to dark red) of BO points.
At 7% noise, the $50^{\rm th}$ percentile sampling has $\mu_{\rm D}({\bf X}^{*})\approx-10$ and even the $75^{\rm th}$ percentile sampling starts to deteriorate. Inspecting how far we are from the GT in ${\bf X}$, we observe a significant deviation with ${\rm\langle IR}({\bf X})\rangle\approx 5$ for the 50th percentile, indicating that less than half the samplings get close to the ${\bf X}_{\rm max}$. Finally, for 10% noise, even the $75^{\rm th}$ percentile does not find the maximum in ${\bf X}$ or $y$. Thus, noise has a clear impact on the convergence rate of BO in heterogeneous type functions, which is one of the key factors to consider when determining the experimental budget for optimization tasks.
Exploration hyperparameters are investigated for both UCB and EI acquisition functions as a function of noise level, based on the performance metric ${\rm\langle IR}({\bf X})\rangle$ (Table  S9 and  S10). Comparing UCB $\beta=1$ and EI $\xi=0$, we find that for noise levels above 3%, the EI acquisition performs better than UCB. The better performance using the EI acquisition function motivated us to compare BO results for different noise levels using EI in Fig. 7. Neither EI nor UCB have been initially designed to be used for noisy objectives,44 and ${\rm\langle IR}({\bf X})\rangle$ are indeed high beyond 10% of noise, indicating most of the LHS samplings have not resulted in a converged BO. Lately, acquisition functions have been designed specifically for noisy objectives, such as Noisy-EI,44, 45 however, these have not yet been integrated into most BO packages widely used in materials science applications. Table  S11 shows the acquisition function and hyperparameter that yields the lowest ${\rm\langle IR}({\bf X})\rangle$ for each noise level, and reveals that for 2% noise, UCB with $\beta=2$ is optimal whereas for 5% and 10%, EI with $\xi$ = 0.05 is optimal, and for 7% noise, EI with $\xi=0$ is optimal. Generally, low $\xi$ values that lead to more exploitation perform clearly better in low-noise scenarios for Ackley, whereas in high-noise scenarios, the choice of $\xi$ does not affect the result drastically. In the case of UCB, the effects of the exploration hyperparameter on the convergence are not as drastic but the low exploration hyperparameter values also perform better in low-noise scenarios. Thus, based on this benchmark, low exploration choices are a robust option for EI and UCB with Ackley-type objectives.
Fig. 7(c) shows the GPR model of the sampling with the $50^{\rm th}$ percentile ${\rm\langle IR}({\bf X})\rangle$ and Fig. 7(d) shows the corresponding parity plots. For 2% noise, the GPR model resembles the GT; for 5%, the peak is slightly degraded; for 7%, the peak is significantly degraded; and finally, for 10%, the landscape is almost flat and the Ackley peak has all but disappeared. A zoom-in 2D heat map is shown in Fig. S15, and the GNV hyperparameters are shown in Fig. S16. Supplementary Movie (5) illustrates the progression of BO learning, showcasing the transformation from the initial to the final iteration of the $50^{\rm th}$ percentile sampling with 5% noise. Overall, we observe good performance at 2% noise with performance degraded as noise increases to almost no BO effectiveness at 10% noise. It is clear from 7(d) that the GPR model misses the peak when noise is above 5%. For 10% noise, the model was only covering $y$ values below -15, significantly lower than the peak. However, for Ackley, these regions constitute almost the entire ($\langle$ 99.99%) GT hypervolume. This fact is important when we compare BO performance for Ackley vs. Hartmann in Fig. 12.
3.2.3 Hartmann
Fig. 8 (a) and (b) show the learning curves of the Hartmann function, using the EI acquisition policy, for noise levels of 2%, 5%, 7%, and 10%. At 2% noise, almost all samplings reach either ${\bf X}_{\rm max}$ or ${\bf X}_{\rm max,2}$; only one LHS sampling results in ${\rm\langle IR}({\bf X})\rangle\approx 0.3$ and $\mu_{\rm D}({\bf X}^{*})<3$ indicating it is neither close to ${\bf X}_{\rm max}$ nor ${\bf X}_{\rm max,2}$. The spread on ${\rm\langle IR}({\bf X})\rangle$ increases with the noise but the 50th percentile sampling still approaches the ${\bf X}_{\rm max}$ even at 10% noise, distinctly different from the Ackley case (Fig. 7).
[…]
font=footnotesize Refer to caption
Figure 8: Learning curve of $||{\bf X}^{*}-{\bf X}_{\rm max}||$ in top row (a) and $\mu_{\rm D}({\bf X}^{*})$ in the second row (b) of the Hartmann function using LP as the batch-picking method and EI as the acquisition function for noise levels of 2% ($\xi$ = 0.005), 5% ($\xi$ = 0.1), 7% ($\xi$ = 0), and 10% ($\xi$ = 0) from left to right. For visualization, (c) 3D representations and (d) parity plots of the 50th percentile BO models for the same noise levels. In the 3D representations, x1 and x2 are projected with 224 learning points, including 24 initial LHS points (red) and 50 iterations (from blue to green) of BO points.
Fig. 8(c) illustrates the GPR model of the 50th percentile sampling and Fig. 8(d) shows the corresponding parity plots. Since the Hartmann function is not symmetric, the projections onto the other pairs of input variables are shown in Fig. S17S20. At low noise (2% and 5%), almost all of the ${\bf X}$ are in the approximate vicinity of ${\bf X}_{\rm max}$. At 7% noise, we observe that a region with $x1\approx 1$ and $x2\in[0.5,1]$ is explored initially but finally ${\bf X}_{\rm max}$ is found. At 10% noise many points are sampled with $x2\in[0,0.5]$ but ${\bf X}_{\rm max}$ is found, and the landscape still visually resembles the Hartmann landscape, albeit with a reduced height at the maximum. The GNV hyperparameters for the noisy cases (Fig. S21) reflect the increasing noise level. BO still converges with a 15% noise level (Fig. S22). The performance of BO degrades with noise with Hartmann test function but, in contrast with Ackley test function, BO remains functional for Hartmann up to 15% noise.
We investigated the optimal exploration hyperparameters for both UCB and EI acquisition functions as a function of noise, based on the performance metric ${\rm\langle IR}({\bf X})\rangle$ (Table  S12 and  S13). Trends similar to the Ackley test function are found for Hartmann, albeit Hartmann test function is less sensitive to the choice. The smallest ${\rm\langle IR}({\bf X})\rangle$-yielding values are used for Fig.8 (EI with $\xi=0.005$, $0.1$, 00, 00 for 2%, 5%, 7%, and 10% noise, respectively). Table  S14 highlights the best-performing acquisition function and hyperparameter for Hartmann test function at each noise level. EI is the best acquisition function for five of the noise conditions whereas UCB is the best for six of the noise conditions with only small differences between each. This indicates that for the Hartmann landscape, UCB and EI have similar BO performance.
Fig. 9 and in Movie (6) (see supplementary) show the full iterative progression of the BO of the Hartmann function with 5% noise using the EI acquisition policy ($\xi$ = 0.1). Initially, the maximum of the GPR landscape is not located near ${\bf X}_{\rm max}$. However, at iteration 32, the optimum has been identified. In comparison, in the noise-free case (Fig.  S8), ${\bf X}_{\rm max}$ and $H({\bf X}_{\rm max})$ are identified after only 5 iterations.
[…]
font=footnotesize Refer to caption
Figure 9: (a)-(f) 3D representations of 50th percentile BO models (EI, $\xi$= 0.1) of the Hartmann test function with 5% noise at different iterations, where x1 and x3 are shown with 224 learning points (including 24 initial LHS points in red). Black cross marks ${\bf X}_{\rm max}$ and red stars marks ${\bf X}_{\rm max,2}$
Fig. 10(a) shows the fraction of LHS samplings where $||{\bf X}^{50}-{\bf X}_{\rm max}||<||{\bf X}^{50}-{\bf X}_{\rm max,2}||$. As noise increases, the fraction of LHS samplings that find ${\bf X}_{\rm max}$ decreases from 75% to 30%. The observation matches with Fig. 8, in which we studied noise up to 10% and observed that the 50th percentile still find ${\bf X}_{\rm max}$. However, further increase in noise results in more than 50% of the LHS samplings ending up closer to ${\bf X}_{\rm max,2}$ compared to ${\bf X}_{\rm max}$. This indicates that for a landscape with almost degenerate maxima, at high noise, the BO is not able to distinguish between the two. This could arise from the area of the convex peak of the second maximum being large compared to the first, as well as the relative locations of the two optima. Whenever the noise level is higher than the difference in the values of the true global optimum and other competing optima, BO has only little evidence for determining which one of the optima is better. However, the ability of BO to find an optimum diminishes more slowly with the increasing noise than the ability to find the global optimum.
[…]
font=footnotesize Refer to caption
Figure 10: For Hartmann function, (a) percentage of data in ${\bf X}_{\rm max}$ as a function of noise levels, and projection of ${{\bf X}^{50}}^{*}$ for all 99 LHS samplings onto the GT 2D heat maps with noise level equal to (b) 0%,(c) 10%, and (d) 20%. Red squares are represent ${{\bf X}^{50}}^{*}$ closer to ${\bf X}_{\rm max}$ and green triangles represent ${{\bf X}^{50}}^{*}$ closer to ${\bf X}_{\rm max,2}$ at the end of 50 iterations.
Locations ${\bf X}^{50}$ are illustrated in Fig. 10(b-d) for 0%, 10%, and 20% noise levels. Without noise, all samplings result in ${\bf X}^{50}$ are clearly near either ${\bf X}_{\rm max}$ or ${\bf X}_{\rm max,2}$. But at 10% noise, a few samplings end up in between the two locations and are attributed to the first or second maximum only through the criterion $||{\bf X}^{50}-{\bf X}_{\rm max}||<||{\bf X}^{50}-{\bf X}_{\rm max,2}||$. The number of ”stray” samplings increases when noise increases to 20%.
3.2.4 Noise Overestimation Problem and Solution
[…]
font=footnotesize Refer to caption
Figure 11: Ackley learning curve of $||{\bf X}^{*}-{\bf X}_{\rm max}||$ (top row) and $\mu_{\rm D}({\bf X}^{*})$ (bottom row) for 10% noise: (a) and (b) of $\rm Max(y_{GT})$ and (c) and (d) of noiseless kernel amplitude. UCB is used as the acquisition policy.
Therefore, we considered other options to incorporate noise in the benchmarking studies. We argue that the kernel amplitudes under noiseless conditions can be taken as the signal level. Thus, a more physical way to set the noise level is setting noise proportionally to the noiseless kernel amplitude rather than $\rm Max(y_{GT})$. This choice stemmed from the observation that the kernel amplitude plays a pivotal role in enhancing the precision of predictions for the objective function during the BO process.
Fig. 11 compares 10% noise case incorporated in two noise frameworks for the Ackley function: proportional to $\rm Max(y_{GT})$ (a-b) and proportional to the noiseless kernel amplitude (c-d). The comparison for the Hartmann function is shown in Fig. S23. The noiseless kernel amplitude is equal to 0.192 and 0.184 for Ackley and Hartmann, respectively. For the kernel amplitude noise framework (Fig. 9(c-d)), all final BO models are well-optimized for the Ackley case. Conversely, for the $\rm Max(y_{GT})$ noise framework (Fig. 11(a-b)), only around 25% of the final BO models exhibit successful optimization, while the remaining fail to achieve satisfactory levels of optimization. The ${\rm\langle IR}({\bf X})\rangle$ value of the $\rm Max(y_{GT})$ is two orders of magnitude larger than that of kernel amplitude noise frameworks for the Ackley case: $31.7\times 10^{-2}$ and $0.38\times 10^{-2}$, reflecting the effect of noise over-estimation on BO outcome.
In the Hartmann case (Fig. S23) with the kernel amplitude noise framework (Fig.. S23(c-d)), again all final BO models converge to ${\bf X}_{\rm max}$ or ${\bf X}_{\rm max,2}$. Conversely, for the $\rm Max(y_{GT})$ noise framework (Fig.. S23(a-b)), only around 40-50% of the final BO models attain successful optimization, while the rest end up in non-optimal regions. However, the noise over-estimation from setting noise level using $\rm Max(y_{GT})$ has a more detrimental effect on heterogeneous problems than non-heterogeneous problems. Adapting the kernel noise framework mitigates this problem and result in good convergence for both functions.
3.2.5 Comparing Noise Effects on Ackley and Hartmann
Fig. 12 (a) and (b) depict ${\rm\langle IR}({\bf X})\rangle$ and ${\rm\langle IR}(y)\rangle$ (normalized by the GT function $\bf X$ and $y$ ranges) as a function of noise level. For Hartmann, ${\rm\langle IR}({\bf X})\rangle\approx 0.2$ without noise and increases to ${\approx 0.35}$ when a small amount of noise is introduced. For Ackley, ${{\rm\langle IR}({\bf X})\rangle}\approx 0$ up to 4% noise after which the ${\rm\langle IR}({\bf X})\rangle$ increases up to 0.4 for 20% noise. The large ${\rm\langle IR}({\bf X})\rangle$ value of Hartmann even without noise is caused by some BO models ending up in ${\bf X}_{\rm max,2}$. The strong dependence of ${\rm\langle IR}({\bf X})\rangle$ for Ackley is the result of noise causing the BO to fail completely to find the optimum. Inspecting Fig. 12b reveals that ${\rm\langle IR}(y)\rangle$ increases rapidly for the Ackley function and reaches 0.5 at 8% noise already due to many of the GPR models not being able to replicate the Ackley peak (see Fig. 7c). For the Hartmann function, even at 20%, $\langle{\rm IR}(y)\rangle\approx 0.1$. Thus, it is more challenging to fit heterogeneous-type objectives.
[…]
font=footnotesize Refer to caption
Figure 12: Noise dependence of (a)normalized $\rm\langle IR(\bf X)\rangle$, (b) normalized $\rm\langle IR(y)\rangle$, and (c) normalized $\rm\langle RMSE\rangle$, and (d) $\rm\langle\sqrt{GNV}\rangle$. All averaged over 99 LHS samplings. For all plots, Ackley test function results are represented by red x’s and Hartmann results are blue +’s. EI is used as the acquisition function and LP is used as batch-picking method.
Fig. 12c shows that $\rm\langle RMSE\rangle$ values (averaged from 99 LHS BO models’ RMSE’s) for both functions increase with noise levels, with the two being approximately the same for $\langle$ 5% noise. At higher noise levels, $\rm\langle RMSE\rangle$ increases much more rapidly for Hartmann compared to Ackley. At first glance, this result is anti-intuitive since the GPR model at 10% noise level does not resemble the Ackley function at all (Fig. 7c), yet the RMSE value is lower than the Hartmann case where the GT landscape is clearly visible (Fig. 8c). This peculiar result is the direct consequence of the Ackley peak occupies only a miniscule fraction of the hypervolume. With the increasing noise, the GPR models completely missed the Ackley optimum region, but instead explore the plateau region where the GT values are small. The small $\rm\langle RMSE\rangle$ between the GT and GPR prediction indicates that BO correctly models the plateau region, and the large ${\rm\langle IR}(y)\rangle$ comes from missing the peak. Finally, the squre root of final model GNV’s averaged over all 99 LHS samplings, $\rm\langle\sqrt{GNV}\rangle$, is shown in Fig. 12d as function of noise. Up to 9% noise, $\rm\langle\sqrt{GNV}\rangle$ values are similar for both the Hartmann and Ackley functions, but beyond 9%, Ackley’s $\rm\langle\sqrt{GNV}\rangle$ becomes slightly lower. From 10% noise onwards, fitting the GPR GNV hyperparameter becomes more challenging for Ackley, while Hartmann’s $\rm\langle\sqrt{GNV}\rangle$ shows a stronger correlation with increasing noise levels, as expected.
4 Conclusion
We investigated high-dimensional Bayesian optimization (BO) and its relevant simulation and design choices in scenarios that involve noise. We have applied batch BO to the 6D Ackley function, representative of a needle-in-a-haystack problem, and the 6D Hartmann function, representative of a problem with a false maximum that are nearly degenerate with the global maximum.
The optimization progression is visualized with learning curves in both $\bf X$ and $y$, evolution curves of hyperparameters, 3D projections of the final Gaussian process regression (GPR) surrogate models and as a function of iterations, and parity plots. With noise, we show that using the optimal posterior mean of the model, $\mu_{\rm D}({\bf X}^{*})$, as the utility function is a more robust way to show BO convergence than using the value of the optimal objective, ${\rm Max}(y)$. These choices on how to track the BO convergence and ensure the proper progression of the optimization gain an increasing importance when BO tasks are performed in high-dimensional search spaces – where the surrogate model landscapes are difficult to visualize comprehensively.
In the absence of noise, BO is able to efficiently find the maximum of the Ackley GT function $({\bf X}_{\rm max})$ whereas for the Hartmann function, 30% of the LHS samplings ended in the false ground truth (GT) maximum $({\bf X}_{\rm max,2})$. We found that in the absence of noise, the UCB acquisition function with exploration hyperparameter $\beta=1$ yielded the best convergence in terms of the lowest instantaneous regret in $\bf X$, ${\rm\langle IR}({\bf X})\rangle$, compared to other values of $\beta$ or acquisition function EI with different values of the exploration parameter $\xi$.
We showed how BO performance is strongly degraded for the Ackley function by noise, with BO model not able to produce the GT maximum when 10% noise is present. For the Hartmann function, we also observe degradation but convergence to the optimum is still maintained up to 15% noise. With increasing noise, a larger fraction of the LHS sampling BO models ends at ${\bf X}_{\rm max,2}$. We inspected ${\rm\langle IR}({\bf X})\rangle$ normalized with respect to the range of the GT function as a function of noise. We found that Hartmann has a large ${\rm\langle IR}({\bf X})\rangle$ in the noiseless case because of the presence of the second maximum but the error increases relatively slowly with the increasing noise level, while ${\rm\langle IR}({\bf X})\rangle$ for Ackley is highly dependent on the noise level, increasing drastically above 4% noise level. These observations show that the specifics of the objective function, i.e., the optimization domain in the experimental case, affect the dependence of the BO performance on the noise, and should thus be evaluated in as much detail as possible when setting up experimental BO.
We also show that simulating noise with respect to $\rm Max(y_{GT})$, as commonly done in the BO literature, is likely a significant over-estimation of the noise that would be seen in experiments, and that simulating noise with respect to the noiseless kernel amplitude of a GPR model fitted on the objective may represent more accurately signal-to-noise ratio in experiments. When using the noise relative to the noiseless kernel amplitude, BO is able to optimize the Ackley function at 10% noise in contrast to the $\rm Max(y_{GT})$ reference level. This highlights the importance of simulating noise in an realistic way prior to experiments, to be able to evaluate the required experimental budget and the feasibility of the optimization correctly.
Overall, we find the convergence of BO deteriorates with increasing noise level, but the behavior depends on the landscape of the optimization domains. Noise makes it harder to locate the global optimum in the Hartmann test function with a false maximum, but does not affect the objective value as much, whereas for the Ackley test function with a sharp peak, missing the location of the peak results in a complete loss in producing the function landscape. This work addresses the challenges of high-dimensional optimization, the importance of optimizing exploration hyperparameter under noisy environment, and introduces visualization techniques and performance metrics to track the progression of BO.
Author Contributions
I.M. - data curation, formal analysis, investigation, software, validation, visualization, writing - original draft, writing - review & editing. A.T. - methodology, validation, writing - review & editing. A.E. and A.S. - software. T.B. - methodology. W.V. - resources, software, supervision, writing - review & editing. J.W.P.H. - conceptualization, funding acquisition, methodology, project administration, resources, supervision, writing - review & editing.
{acknowledgement}
We thank M. Lee, R. Garnett, B. Das, K. Brown, and K. Snapp for helpful discussions. We acknowledge the Texas Advanced Computing Center (TACC) at the University of Texas at Austin for providing the high-performance computing resources that have contributed to the research results reported within this paper. URL: http://www.tacc.utexas.edu.
{suppinfo}
The supplementary materials accompanying this paper include
1. A PDF file with mathematical description and complete 3D visualization of the test functions, performance metrics for different exploration hyperparameter values for both acquisition functions and test functions, plots of kernel hyperparametr evolution during the optimization for both test functions, Ackley results with optimum location not at the center of the hypervolume, 3D visualization of Hartmann optimization results, comparisons between UCB and EI learning results for both test functions, comparisons of local penalization, Kriging believer, and constant liar for both acquisition functions and test functions, optimal acquisition function and hyperparameter value for noise levels from 0 to 20 % for both functions using ${\rm\langle IR}({\bf X})\rangle$ as the metric, and additional results for different noise levels.
2. Six animated videos illustrating the optimization process where the iteration progress is represented through the time evolution: (1) the $\rm 50^{th}$ percentile GPR model projected to x1-x2 plane for the noiseless Ackley function; (2) the $\rm 50^{th}$ percentile GPR model projected to x1-x3 plane for the noiseless Hartmann function; (3) the $\rm 25^{th}$ percentile GPT model (left) and (4) $\rm 75^{th}$ percentile GPR model projected to x1-x3 plane for the noiseless Hartmann function; (5) the $\rm 50^{th}$ percentile GPR model projected to x1-x2 plane for the Ackley function with 5 noise; (6) the $\rm 50^{th}$ percentile GPR model projected to x1-x3 plane for the Hartmann function with 5 noise.
Funding
This work is supported by National Science Foundation CMMI-2109554. A.T. acknowledges funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 01059891. J.W.P.H. and T.B. acknowledge the support from the Simons Foundation Pivot Fellowship. J.W.P.H. acknowledges the support of the Texas Instruments Distinguished Chair in Nanoelectronics.
Data availability
Python codes and data reported in the main text and supplementary materials are available on GitHub: https://github.com/UTD-Hsu-Lab/Noisy-BO. During the review, reviewers can access the same information via https://utdallas.box.com/s/kb3u37odpok0bfgs93ijkb3mnmdc0h6n.